A 20-inch wire will be cut into two parts. One part is bent to form a square and the other
into a circle. If the wire is cut x units from one end, express the sum of the areas of the figures formed
in terms of x.
1
Expert's answer
2020-09-24T18:10:42-0400

A wire of 20 inches is cut x inches from one end

First part forms square of perimeter x inches

The other part forms circle of circumference (20 - x) inches as shown in the figure below


For the Square

"P = 4L"

"x = 4L"

"L = \frac{1}{4}x"

"Therefore, A = L ^2"

"A = (\frac{1}{4}x)^2"

Area of square in terms of x "= \frac{1}{16}x^2"


For the Circle

"Circumference, C = 2\pi r"


"20 - x = 2\pi r"


"r = \frac{20 - x}{2\pi}"


"Area, A = \pi r^2"


"= \frac{22}{7} × (\frac{20 - x}{2\pi} ) ^2"



"= \frac{22}{7} × \frac{(20 - x)^2× 49 }{4 × 484}"


"=\frac{1078 (20 - x)^2 }{13552}"

"=\frac{1078 (400 - 40x + x^2) }{13552}"

"=\frac{7 (400 - 40x + x^2) }{88}"

"= \frac{2800 - 280x + 7x^2}{88}"

Area of circle in terms of x "= \frac{350}{11} - \frac{35}{11} x + \frac{7}{88}x^2"

Sum of area of the figures formed "=\frac{1}{16}x^2 + \frac{7}{88}x^2 - \frac{35}{11} x +\frac{350}{11}"

"= (\frac{25}{176}x^2 - \frac{35}{11} x +\frac{350}{11} ) Sq. Inches"